As a store is 5 times as long as it is wide. If the perimeter of the store is 240 feet. How wide is the store?

Answer by macston(5194)

width=W; Length=L=5W
P=Perimeter=2(L+W)=240 ft
2(L+W)=240 ft Substitute for L
2(5W+W)=240 ft
12W=240 ft divide each side by 12
W=20 feet The width of the store is 20 feet
CHECK
L=5W=5(20 ft)=100 ft
P=2(L+W)
240 ft=2(100ft+20ft)
240 ft=2(120 ft)
240 ft=240 ft